OpenStudy (anonymous):

What is the LCM of 14 and 35? A. 140 B. 70 C. 56 D. 7

iYuko (iyuko):

What is the least common multiple of your numbers?

iYuko (iyuko):

Find the multiples of 14

OpenStudy (anonymous):

Yes, do as @iYuko has said and find the multiples of 17 and then 35. :)

OpenStudy (anonymous):

*14

OpenStudy (anonymous):

B is the correct answer

OpenStudy (anonymous):

how @geerky42

OpenStudy (anonymous):

Do Not Give Out Direct Answers. @sourav_aich

OpenStudy (anonymous):

do i multiply 14 times a number

iYuko (iyuko):

2*7 is 14

OpenStudy (anonymous):

No, no. To find the answer find some multiples of 14.

OpenStudy (anonymous):

The only multiples of 14 are {1, 2, 7, 14}. Can you name some multiples of 35? @BrieiaKlarc

OpenStudy (anonymous):

lol those are factors

OpenStudy (anonymous):

answer B i.e 70 is the correct answer @BlossomCake

OpenStudy (anonymous):

how do you know that

iYuko (iyuko):

@sourav_aich DIRECT ANSWERS ARE VIOLATING OPENSTUDY CODE OF CONDUCT. PLEASE REFRAIN FROM DOING IT.

OpenStudy (anonymous):

B???

OpenStudy (anonymous):

LCM means least common multiple so 70 is the least common multiple of 14 and 35

OpenStudy (anonymous):

I am sorry for misunderstanding the problem. Yes, to find the answer find some common multiples between 14 and 35.

OpenStudy (anonymous):

is B the correct answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@BrieiaKlarc List some multiples of 14.

iYuko (iyuko):

Solve it. Show proof that that is the right answer. Because 70 could be wrong.

OpenStudy (anonymous):

14 · 0 = 0 14 · 1 = 14 14 · 2 = 28 14 · 3 = 42 14 · 4 = 56 14 · 5 = 70 14 · 6 = 84 14 · 7 = 98 14 · 8 = 112 14 · 9 = 126

iYuko (iyuko):

Now do 35. :)

OpenStudy (anonymous):

Multiples of 14: {14, 28, 42, 56, 70, 84, 98}

OpenStudy (anonymous):

Good, good. Do the same with 35.

OpenStudy (anonymous):

70 105 140 175 210 245 280 315 350

OpenStudy (anonymous):

LCM of two numbers 'a' and 'b' is the smallest positive number that is divisible by both 'a' and 'b' so 70 is the least number that is divisible by both 14 and 35

OpenStudy (anonymous):

can you find any positive number less than 70 that is divisible by both 14 and 35 ?

OpenStudy (anonymous):

B is the answer?? @iYuko @BlossomCake

OpenStudy (anonymous):

Yes, B. 70 is the answer. :)

OpenStudy (anonymous):

ok